# Engineering morphogenesis of cell clusters with differentiable programming

### v2

@@ Abstract
Understanding the rules underlying organismal development is a major unsolved problem in biology. Each cell in a developing organism responds to signals in its local environment by dividing, excreting, consuming, or reorganizing, yet how these individual actions coordinate over a macroscopic number of cells to grow complex structures with exquisite functionality is unknown. Here we use recent advances in automatic differentiation to discover local interaction rules and genetic networks that yield emergent, systems-level characteristics in a model of development. We consider a growing tissue with cellular interactions mediated by morphogen diffusion, differential cell adhesion and mechanical stress. Each cell has an internal genetic network that is used to make decisions based on the cell’s local environment. We show that one can simultaneously learn parameters governing the cell interactions and the genetic network for complex developmental scenarios, including the symmetry breaking of an embryo from an initial cell by creation of emergent chemical gradients, homogenization of growth via mechanical stress, programmed growth into a pre-specified shape, and the ability to repair from damage. When combined with recent experimental advances measuring spatio-temporal dynamics and gene expression of cells in a growing tissue, the methodology outlined here offers a promising path to unravelling the cellular bases of development.
@@
Inverse design tasks are usually translated into optimization problems, wherein the objective is to minimize the discrepancy between a property of the simulated system state and its desired value. While sophisticated computational models of cell behavior exist [13, 14], their applicability to reverse engineering remains limited, particularly in high-dimensional parameter spaces where parameter sweeps and evolutionary searches become inefficient. Automatic differentiation (AD) algorithms, on the other hand, enable efficient calculation of sensitivities of observables with respect to input parameters. While originally developed for training large neural networks [15, 16], the same technical infrastructure can be used to differentiate through simulations of physical processes, offering gradient-based optimization as a viable alternative to tackle hard inverse problems of various scientific interest. Successful examples include the design of self-assembling materials [17], the establishment of data-driven discretizations for partial differential equations [18], nonequilibrium statistical mechanics [19] and the design of quantum devices [20]. Neural Cellular Automata [21] represented an initial demonstration of how these methods could be used to engineer collective behaviors inspired by biology in lattice models. Hiscock, instead, showed that AD could be used to successfully learn meaningful couplings in ODE models of genetic networks [22].
Here, we consider a model of a tissue composed of cells, where each cell can divide, grow, exert stress, as well as excrete and sense morphogens. Cellular decisions are made in response to the local environment, via an internal genetic network (Fig. 1). Starting from random initial parameters, we demonstrate that we can learn the parameters governing both cellular and genetic interactions so that the cluster of cells undergoes non-trivial developmental scenarios, ranging from symmetry breaking of an embryo from a progenitor cell, to the creation of emergent chemical gradients, to homogenization of growth via chemical signalling and mechanical stress response, to programmed growth into a pre-specified shape, which is able to repair itself from damage. The framework allows for the probing of the relationship between the capabilities of individual cells and the emergent properties of the resulting tissues.
@@
Figure 1: Modeling and Optimization Framework. (a) Model overview. Decision-making circuits in cells sense local chemical and mechanical inputs to regulate secretion, division probability, and adhesion. The environment responds to cellular outputs by handling chemical diffusion, cell division, and mechanical relaxation. (b) Input signals are integrated and elaborated in each cell by internal gene regulatory networks. Green gene nodes sense inputs, blue genes are intermediates and red genes can modify cellular actions. Gene expression levels in each node are maintained from one simulation step to the next, providing memory to the system. (c) Gradient calculation and optimization. Following the REINFORCE approach, the log probability gradient of each time step is weighted by action rewards to estimate the average gradient and parameters are updated through gradient descent. Gradients flow through the whole differentiable simulation, accounting for environment contributions.
@@
We model tissue growth using soft spheres interacting through a Morse potential [23]. We mostly consider two-dimensional systems since three-dimensional systems rapidly become too computationally intensive to optimize with our current implementation. Our methods are nonetheless applicable in principle to any physical model of cell interactions and we show some simple experiments with 3D systems. In our model, cells grow at a fixed rate until reaching a maximum radius, leading to continuous mechanical reorganization. Cells secrete $M$ morphogens, measure local concentrations and spatial gradients of these chemicals, and sense mechanical stress. This locally sensed information guides decisions on chemical secretion modulation and division probability (Fig. 1a). Cell division events are stochastic, producing two daughter cells, each with half the volume of a fully grown mother cell. Decision-making circuits are modeled as regulatory networks comprising $N$ interacting genes with trainable excitatory or inhibitory couplings (Fig. 1b). The regulatory network processes the information sensed by individual cells to output rates of division, chemical secretion, etc. In most examples, the regulatory network weights are shared by all cells, though we also explore one scenario where two cell types have partially different regulatory networks. Further mathematical detail can be found in the Supplementary Information.
@@
We employ score-based methods such as REINFORCE [27] to address the intrinsic stochasticity of growth dynamics, which implies that the loss function is not necessarily differentiable (Fig. 1c). In the REINFORCE update, rewards and penalties are assigned to each action taken (i.e. each cell division) to provide the weighting signal for the gradient of the trajectory probability. This weighted gradient is then used to nudge the weights of the gene network in the right direction. In practice, the system gradually learns which division events are the most favorable and increases their probability for the next simulation. To guide the reader possibly unfamiliar with these concepts, we also include a set of in-depth tutorial notebooks in the GitHub repository [28] and a longer theoretical explanation in the Supplementary Information. Since the dynamics of the environment, including chemical secretion and diffusion, depend partly on the same parameters (the gene network couplings) that cells use to decide whether to divide or not (their “policy”), it is essential to evaluate the gradient of the environment dynamics with respect to the gene network parameters. This necessity makes automatic differentiation particularly crucial in this context. Once the necessary gradients are calculated, optimization is performed using well-established gradient descent schemes, such as Adam [29]. The learned networks are then further simplified by removing all edges with weights smaller than an experiment-dependent threshold, such that the functional backbone of pruned network be highlighted while preserving the optimized properties.
Figure 2: Chemical Regulation of Cellular Adhesion. Learned chemical regulation of cellular adhesion guides the assembly of different spatial cluster arrangements. The top part of the panels show the training loss curves for three different optimization targets. The insets show the final state of a simulation after the learning, with colors denoting different cell types. The bottom sections show the learned and pruned gene network; Cad1, Cad2, and Cad3 are the homotypic cadherins of the three cell types, and Het is the heterotypic cadherin expressed in all cell types. (a) Lattice-like structure. Cells are arranged to minimize the sum of pairwise distances between heterotypic cells. (b) Lobed structure. Cells minimize the sum of pairwise distances between homotypic cells. (c) Core-shell structure. Cells are arranged to minimize the distance of cells from a predefined radius, dependent on their type. The final configuration is here shown as a cross-section to highlight the internal structure.
Chemical Regulation of Cellular Adhesion
The differential adhesion hypothesis posits that populations of cells can adopt specific morphologies based on differences in adhesive strength [30]. During morphogenesis, a cadherin code regulates the spatial organization of developing tissue [31, 32]. Engineered cellular circuits can use gene regulatory networks to dynamically modulate adhesion and program self-organization [6]. Here, we show that gene interactions can be learned that use chemical signaling to modulate homotypic and heterotypic adhesion to achieve a target spatial pattern in a population with different cell types.
In our model, a population of three cell types is subject to Brownian motion and interacts via a Morse pairwise potential. The well depth of the potential between a pair of cells is dictated by the concentration of homotypic and heterotypic “cadherins” in each cell. Each cell type produces its respective homotypic cadherin and a shared heterotypic cadherin, regulated by chemical signaling. Each cell type utilizes the same gene network, but can only secrete chemicals and express cadherins of its own type. To achieve a desired spatial organization, we construct three loss functions based on pairwise distances between cells. For a lattice-like structure, we minimize the sum of pairwise distances between cells of different types (Fig 2a). To organize the structure into three lobes, we minimize the sum of pairwise distances between cells of the same type (Fig 2b). To achieve a core-shell structure, we minimize the distance of each cell to a pre-specified radius, based on the cell type (Fig 2c).
The optimization process identifies effective solutions by allowing cells to observe their own type and regulate the pairwise potential between cell types. For a lattice-like structure, cells up-regulate their heterotypic cadherin and down-regulate their homotypic cadherin. Conversely, to create a lobed structure, they up-regulate their homotypic cadherin and down-regulate their heterotypic cadherin. These patterns of homo- and heterotypic adhesion are consistent with observations from cell culture sorting experiments [31]. Achieving core-shell structures requires a more complex regulatory network (Fig 2c).
@@
Understanding how to program the growth of complex shapes can aid in our comprehension of developmental body plans driven by chemical patterning. Rationally designed synthetic gene circuits have already been shown to be able to drive the formation of target multicellular structures found in morphogenesis [6, 33]. We investigate here how to use our framework to discover new decision-making circuits that utilize local chemical interactions to grow pre-specified shapes. These rules are, as above, formulated as gene regulatory circuits. We focus in particular on two fundamental shapes commonly found in developmental biology: elongation and branching. The symmetry-breaking elongation process is the first geometric-defining process to happen in many developmental programs and is vital, for example, for establishing the antero-posterior axis in animals [34]. Branching is possibly the next most ubiquitous process in the development of multicellular organisms and branched structures are found from plant roots to blood vessels and mammary epithelial cultures [35].
We show below that gene regulations that achieve these final structures can be learned in our model. Notably, we also show that mechanisms optimized solely to achieve a specific final structure exhibit emergent properties that they were not explicitly trained to achieve. In particular, gene circuits learn to leverage, modify and generate chemical gradients and they display basic regenerative capabilities.
Figure 3: Elongation. Maximization of the squared sum of the z-coordinates of cells, resulting in vertical elongation in the optimized cluster. (a) Loss curve for one successful optimization run. On the left the final configuration of a simulation with randomly initialized parameters, on the right the final state after learning. (b) Visualization of the optimized and pruned gene network, displaying activation and inhibition interactions that regulate cell behavior during elongation. (c) Two examples of emerging chemical gradients that are used to guide growth.
Elongation
We consider a three dimensional model of dividing cells that we optimize to minimize the squared sum of the z-coordinates of cells, thus requiring the cluster to elongate vertically (Fig. 3a). Local chemical concentrations, local estimated chemical gradients, cell size, and division propensities are passed as inputs into the cells’ gene network (Fig. 3b), that infer positional information through local physical communication. Simulations start from a single cell and grow to a total of 120 cells.
In order to globally assume an elongated shape, cells learn to create emergent chemical gradients or control existing ones, even though this task is not included in the optimization objective (Fig 3c). The cells learn to divide primarily at one or both ends of the growing cluster. Since the cluster edge itself moves as a consequence of new cells being added, the gene networks must implement a non-trivial control mechanism to keep the cluster growing.
Figure 4: Branching. Optimizing growth in a V-shaped reward landscape results in a learned branched structure of the final cluster. (a) Loss curve for one successful optimization run. On the left, the final configuration of a simulation with randomly initialized parameters. On the right, the final state after learning with a visualization of the reward mask (red areas are associated to penalties, white ones with rewards). (b) Two examples of emerging chemical gradients that are used to guide growth. (c) Final optimized state after 129 cell divisions, the same number used for the training (top) and with 500 extra cell divisions. Note the spontaneous branching on the main arms. (d) Loss of the final state after removal, at random, of a fraction of cells. The red dashed line is the loss associated to the final simulation state before cell removals.
Branching
We simulate the growth of a cluster from a single cell to 130 cells, restricting ourselves to the 2-dimensional case for computational simplicity. With a similar spirit to diffusion-based geometry [35], we simulate permeable diffusive boundaries between the cell cluster and the surrounding implicit medium by connecting boundary cells to a ghost chemical sink node. Gene networks have access to the local chemical concentrations, estimates of local chemical gradients, their own cell radius and division propensity. They do not have access to their own absolute spatial coordinates.
To learn more complex branched structures, we construct a position-based shape loss. At each time point, cell divisions that place cells within the region designating the desired shape are assigned a reward, the others a penalty. In this case, we impose a V-shaped spatial reward structure (Fig. 4a on the right) augmented with an asymmetry penalty on the x-coordinate, in order to promote the growth of both arms at the same time. Also in this case, a variety of chemical patterns emerge in the learned simulation (Fig 4b), and the cluster learns to divide cells primarily at the tips of the branching structure. Clusters visually acquire a branched shape, even when simulated for 500 cell divisions more than optimized for. While shape errors clearly emerge, they do mostly in the form of smaller branches that begin to sprout from the main branched structure (Fig 4c). Furthermore, the learned network displays emergent regenerative capabilities. After reaching the final branched configuration, we remove a fraction of cells chosen at random and we allow the cluster grow back to its original size. As shown in Fig 4d, eliminating up to 60% of cells still results in the regeneration of a branched structure. Even more interestingly, it appears that randomly removing cell actually promotes error-correction in the regrowth, reaching even slightly better performances in the training loss.
Figure 5: Mechanical Mechanisms for Homogeneous Growth. (a) Schematic of setup of optimization problem for learning uniform growth and definition of loss function. (b) Visualization of the final configuration of the cluster from an optimized and unoptimized simulation. The unoptimized simulation does not regulate division propensities based on local inputs and thus, growth reflects the growth factor gradient. Cells are colored by the division propensity. (c) Visualization of trained gene network. (d) Gene network learned with a power law morphogen field readily works on three other kinds of morphogen fields: an exponential field, a field with two peaks instead of one, and a field with an off center peak. Losses are shown for each field type, averaged over 100 simulations. (e) Division rate divided by the morphogen field concentration is plotted for each kind of simulation from (d). This shows how cells regulate division probabilities without the influence of the field; results show that the way cells regulate division adapts to the kind of field due to the learned stress-based mechanism.
@@
Mechanical signaling has long been recognized as an important component of morphogenesis. In individual cells, mechanosensation influences gene expression through several identified biochemical pathways that can transduce forces [36]. On a tissue-level scale, mechanical signals can communicate local instructions to build a global shape [36, 37, 38, 39]. In tissues, mechanical interactions can act on larger length scales than biochemical signals due to the propagation of tension through connected cells [39]. The buildup of local mechanical stress, additionally, can influence global growth of a developing structure. As an example, mechanical feedback has been posited as a mechanism for growth arrest in the wing imaginal disc of Drosophila [40, 41]. Here, we explore these ideas by optimizing over a model inspired by Hufnagel et al. to learn mechanisms that couple mechanics to growth.
In our model, a radial growth factor gradient scales a division propensity calculated by gene networks in each cell, such that cells closer to the origin have a higher probability of division than cells at the edges (Fig. 5a). The growth factor gradient follows a power law and is fixed with respect to the simulation. We run the simulation for 200 cell division steps starting from one cell, with each cell modulating its secretion of two morphogens and its division propensity. We seek local growth rules that achieve homogeneous growth (homogeneous division propensities) throughout the cluster despite the nonuniform effect of the growth factor. The gene network that calculates the propensity of division does not see the growth factor as an input; therefore, cells have to infer this information by sensing stress or communicating through morphogens. The learned gene network can sense morphogen concentrations, morphogen gradients, and stress as cellular inputs. We optimize couplings between genes that sense these inputs, four internal genes, and three output genes that modulate secretion rates of the two morphogens as well as the division propensity. The optimization minimizes the coefficient of variation of division propensities at the end of the simulation:

### v3

@@ Abstract
Understanding the rules underlying organismal development is a major unsolved problem in biology. Each cell in a developing organism responds to signals in its local environment by dividing, excreting, consuming, or reorganizing, yet how these individual actions coordinate over a macroscopic number of cells to grow complex structures with exquisite functionality is unknown. Here we use recent advances in automatic differentiation to discover local interaction rules and genetic networks that yield emergent, systems-level characteristics in a model of development. We consider a growing tissue with cellular interactions mediated by morphogen diffusion, cell adhesion and mechanical stress. Each cell has an internal genetic network that is used to make decisions based on the cell’s local environment. We show that one can learn the parameters governing cell interactions in the form of interpretable genetic networks for complex developmental scenarios, including directed axial elongation, cell type homeostasis via chemical signaling and homogenization of growth via mechanical stress. When combined with recent experimental advances measuring spatio-temporal dynamics and gene expression of cells in a growing tissue, the methodology outlined here offers a promising path to unraveling the cellular bases of development.
@@
Inverse design tasks are usually translated into optimization problems, wherein the objective is to minimize the discrepancy between a property of the simulated system state and its desired value. Pioneering research showcased the potential of using evolutionary algorithms to design genetic networks with specified functions [13]and and gaining insights into collective developmental cellular behaviors like segmentation [14]. While sophisticated computational physical models of cell behavior do exist [15, 16], their applicability to reverse engineering remains limited, particularly in high-dimensional parameter spaces where parameter sweeps and evolutionary searches become inefficient. Automatic differentiation (AD) algorithms, on the other hand, enable efficient calculation of sensitivities of observables with respect to input parameters. While originally developed for training large neural networks [17, 18], the same technical infrastructure can be used to differentiate through simulations of physical processes, offering gradient-based optimization as a viable alternative to tackle hard inverse problems of various scientific interest. Successful examples include the design of self-assembling materials [19], the establishment of data-driven discretizations for partial differential equations [20], nonequilibrium statistical mechanics [21] and the design of quantum devices [22]. Neural Cellular Automata [23] represented an initial demonstration of how these methods could be used to engineer collective behaviors inspired by biology in lattice models. Hiscock, instead, showed that AD could be used to successfully learn meaningful couplings in ODE models of genetic networks [24].
In this study, we consider a tissue model composed of cells capable of division, growth, mechanical stress sensing, and morphogen excretion and detection. Cellular decisions are driven by responses to the local environment through an internal genetic network (Fig. 1a). We demonstrate that, starting from random or uninformative initializations, we can optimize the parameters governing genetic interactions to achieve non-trivial developmental scenarios in cell clusters. Our examples draw inspiration from various biological processes: axial elongation during limb bud outgrowth, maintenance of cell type homeostasis through chemical signaling in macrophage-fibroblast co-cultures, and growth homogenization via chemical signaling and mechanical stress response in the Drosophila wing disc. Our framework enables the probing of the complex relationship between individual capabilities of proliferating cells and emergent properties of the resulting tissues.
@@
Figure 1: Modeling and Optimization Framework. (a) Model overview. Decision-making circuits in cells sense local chemical and mechanical inputs to regulate secretion, division probability, and adhesion. The environment responds to cellular outputs by handling chemical diffusion, cell division, and mechanical relaxation. (b) Input signals and internal state are integrated and elaborated in each cell by internal gene regulatory networks. Green gene nodes sense inputs, blue genes are intermediates and red genes can modify cellular actions. (c) Released chemical factors diffuse in the environment following steady-state diffusion. (d) Mechanical interactions between cells are modeled with a pairwise Morse potential. The repulsive region accounts for volume exclusion forces and the attractive region models cell adhesion. (e) Gradient calculation and optimization. Following the REINFORCE approach, the log probability gradient of each time step is weighted by action rewards to estimate the average gradient. Parameters are updated through gradient ascent to maximize the expected reward.
@@
We model proliferating cells in a tissue with adhesive soft spheres, demonstrating examples of both two- and three-dimensional systems. While we present specific examples in this paper, our methods are potentially applicable to various physical models of cell growth and interaction. In our model, cells grow at a fixed rate until reaching a maximum radius, leading to continuous mechanical reorganization. Cells secrete morphogens, measure local concentrations of these chemicals, and sense mechanical stress. This locally sensed information guides decisions on chemical secretion modulation and division probability (Fig. 1a). Cell division events are stochastic and produce two daughter cells, each with half the volume of a fully grown mother cell; the division plane is chosen uniformly at random. Decision-making circuits are modeled as regulatory networks comprising interacting genes with trainable excitatory or inhibitory couplings (Fig. 1b). The regulatory network processes the information sensed by individual cells to output division propensity, rates of chemical secretion, etc. In most examples, the regulatory network weights are shared by all cells, though we also explore one scenario where two cell types have separate regulatory networks. Secreted chemical factors allow information sharing by diffusing in the environment (Fig. 1c). Cells mechanically interact through a Morse potential (Fig. 1d), with a repulsive part for excluded volume interactions and an attractive part for cell-cell adhesion [25]. Simulations are conducted for a specific number of cell divisions, so the final state cannot in general be considered a steady state. Further mathematical detail can be found in the Materials and Methods section and in the Supplementary Information.
@@
We employ score-based methods such as REINFORCE [29] to address the intrinsic stochasticity of proliferation dynamics, which implies that the loss function is not directly differentiable (Fig. 1e). In the REINFORCE update, rewards and penalties are assigned to each action taken (i.e. each cell division) to provide the weighting signal for the gradient of the trajectory probability. This weighted gradient is then used to nudge the weights of the gene network in the right direction. In practice, the system gradually learns which division events are the most favorable and increases their probability for the next simulation. To guide the reader possibly unfamiliar with these concepts, we also include a set of in-depth tutorial notebooks in the GitHub repository [30] and a longer theoretical explanation in the Supplementary Information. Since the dynamics of the environment, including chemical secretion and diffusion, depend partly on the same parameters (the gene network couplings) that cells use to decide whether to divide or not (their “policy”), it is essential to evaluate the gradient of the environment dynamics with respect to the gene network parameters. This necessity makes automatic differentiation particularly crucial in this context. Once the necessary gradients are calculated, optimization is performed using well-established gradient descent schemes, such as Adam [31]. The learned networks are then further simplified by removing all edges with weights smaller than an experiment-dependent threshold, such that the functional backbone of pruned network be highlighted while preserving the system properties we optimized for.
Figure 2: Elongation. Maximization of the squared sum of the x-coordinates of cells, resulting in horizontal elongation in the optimized cluster. (a) On the left, the final configuration of a simulation with randomly initialized parameters; on the right, the final simulation state after learning. Source cells (in red) secrete the growth factor and cannot divide. Proliferating cells (in gray) sense the growth factor and divide in response to it. (b) The learned gene network. The receptor gene is activated only by the presence of the external chemical factor, which results in repression of the division propensity. Edge boldness is correlated to connection strength. (c) Chemical gradient created by source cells along the cluster x-coordinate. In the inset, the same gradient visualized in physical space. (d) Division propensity distribution at the end of the simulation along the x-axis, highlighting the concentration of dividing cells at the tip.
@@
Understanding how to program the growth of complex shapes can aid in our comprehension of developmental body plans driven by chemical patterning. Rationally designed synthetic gene circuits have been shown to be able to drive the formation of target multicellular structures found in morphogenesis [6, 32]. We investigate here how to use our framework to discover decision-making circuits that utilize local chemical interactions to grow target shapes. As an example, we focus in this section on symmetry-breaking axial elongation. This process is the first geometric-defining process to happen in many developmental programs and is vital for establishing, for instance, the antero-posterior axis in animals [33]. We learn a gene circuit that enables cells to chemically interact with a non-proliferating population of ”source cells” to elongate along a given axis. Similar problems have recently been explored by engineering, for example, contact-based cell communication instead of chemical communication [34, 35]. We consider a three-dimensional model of dividing cells, that we optimize to minimize the squared sum of the x-coordinates of cells, thus requiring the cluster to elongate horizontally (Fig. 2a). The system is composed of two different cell types. One secretes a diffusible factor (source cells, in red in Fig. 2a), but cannot proliferate. The other cell type (in gray in Fig. 2a) can only sense the factor concentration and modulate its division propensity accordingly. Local chemical concentration is the input to the proliferating cells’ gene network, allowing them to infer positional information through local physical cues.
The learned mechanism (Fig. 2b) drives a dynamic process reminiscent of limb bud outgrowth in animal development. Secretion of the chemical factor by source cells results in a steady-state chemical gradient along the cell cluster (Fig. 2c). Due to the network’s input-to-output strong inhibitory link, regions of higher chemical concentration, proximal to the source cells, experience a decrease in the division propensity of proliferating cells. Conversely, areas with lower chemical concentration experience less inhibition, leading to a sustained higher division propensity located in the distal region. In the initial spherical cluster, source cells are localized on the left and proliferating cells on the right, the latter cells possessing a non-zero initial division propensity. Proliferating cells closer to the source will experience a diminishing division propensity over time due to the presence of the chemical factor. As these cells divide, they preferentially expand the cluster in directions of lower chemical concentration, effectively moving away from the source and further reducing the chemical concentration at the extending tip. This creates a self-reinforcing loop: cell division events are concentrated at the distal tip where the chemical signal is weakest, ensuring directional growth and elongation away from the source of the secreted factor (Fig. 2d). Crucially, the weak inhibitory feedback from division propensity back to chemical input acts as a contrast enhancer, sharpening the division response to the chemical gradient and producing a ”winner-takes-all“ kind of behavior. In areas of high chemical concentration, the strong inhibitory forward connection suppresses division, thus diminishing the input suppression by the weaker feedback connection and leading to even stronger division suppression. Conversely, in regions of low chemical concentration where direct inhibition is less prominent, division propensity increases and the weak feedback becomes functionally significant, further reducing the perceived chemical input. This effect amplifies the division propensity output specifically in these low-concentration regions. The positive self-loop on division propensity further stabilizes and amplify this directional growth by sustaining high division rates in the low-concentration tip regions, and working in synergy with the contrast-enhancing feedback to create a robust and spatially well-defined elongation.
While significantly simplified compared to biochemical networks governing directional growth, the learned architecture functionally recapitulates the core logic of some developmental processes. Like limb bud outgrowth, it achieves directional elongation through localized growth at the tip [36]. Unlike the growth-promoting signaling factors (like FGF) secreted by the Apical Ectodermal Ridge, though, it employs an inhibitory regulatory signal originating from the base to restrict proximal proliferation and promote growth at the distal end. This regulation is learned consistently in all our optimization runs. The lack of alternative mechanisms (e.g. maintaining a proliferating zone close to the source cells) might be due to different reasons, like the particular mechanical properties chosen to simulate the system, the absence of external geometric constraints or the specific reward function chosen.
Figure 3: Chemical Regulation of Homeostasis. (a) Two cell types each exclusively secrete a chemical factor that can be sensed by both cell types. (b) Visualization of the learned gene networks. Interactions in black are present in every learned network. Colored interactions are single variants that are present only in a subset of learned solutions, with the reported frequency. This gives three type 1 network variants and two type 2 variants, totaling to 6 possible overall regulatory combinations. Edge boldness is correlated to connection strength. (c) An initially imbalanced state (in the middle) grows into an equally unbalanced final state with untrained network parameters (left) and into a balanced state with learned parameters (right) (d) Average over 100 realizations of final proportion of red cells at the end of simulations of increasing lengths. (e) Average over 100 realizations of final proportion of red cells at the end of simulations with increasing proportions of red cells in the starting state. In both plots the shaded area corresponds to average plus or minus one standard deviation.
Chemical Regulation of Tissue Homeostasis
Maintaining stable ratios of different cell populations is crucial for tissue function, and disruptions of this balance can lead to disease [37]. To maintain these stable proportions, cellular circuits must prevent uncontrolled growth of any single cell type. Cell communication mediated by diffusible factors constitutes a key mechanism for achieving tissue homeostasis and regulating tissue composition.
One prominent example is the interplay between osteoblasts and osteoclasts in bone homeostasis [38]. Bone-forming osteoblasts primarily secrete RANKL, which promotes osteoclast differentiation and activation. Conversely, osteoclasts secrete their own factors (like S1P) and release bone matrix factors (e.g. TGF-$\beta$) during bone resorption, influencing osteoblast abundance. This reciprocal signaling establishes a feedback loop that maintains the balance between bone formation and resorption. Similar phenomena are also observed in vitro. For instance, co-cultures of macrophages and fibroblasts can spontaneously reach stable ratios, even from highly imbalanced starting conditions [39]. Both cell types secrete growth factors that selectively stimulate the growth of the other: macrophages primarily produce PDGF and fibroblasts secrete Csf1, while other diffusible factors contribute to self-regulation of growth. Cell-to-cell chemical communication via diffusible factors has also been shown to be crucial for regulating homeostasis during cell differentiation [40, 41].
Inspired by the macrophage-fibroblast system in particular, we investigated chemical mechanisms capable of maintaining a balanced 50/50 ratio between two cell types in a growing cluster. In this 2D model, two cell types interact chemically, each secreting a distinct chemical factor (Fig. 3a). Cell behavior is governed by cell-type-specific genetic networks, learned through optimization, which regulate division propensity and secretion of diffusible factors (Fig. 3b). As shown in Fig 3c (center), simulations begin with an imbalanced initial state, comprising $n_{r}=4$ red cells (cell type 1) and $n_{b}=16$ blue cells (cell type 2). The untrained network is initialized to yield uniform division propensities: after 100 cell divisions, the initial imbalance is found again in the final state (Fig 3c, left). The learned mechanisms, instead, successfully yield balanced cell type populations at the end of the simulation. Gene couplings are optimized to minimize the number difference between the two cell types during growth (Fig 3c, right).
$\mathcal{L}_{t}=|n_{r}(t)-n_{b}(t)|$
is the loss at time $t$, where $n_{r}$ and $n_{b}$ are the number of red and blue cells respectively, and $n_{r}+n_{b}=N$ is the total cell number.
The learned networks for each cell type are shown in Fig. 3b. Both cell types sense both chemicals, but each modulates the secretion of its own factor only. We focus first on the connections in black in the figure, representing the core ”backbone” network consistently learned across optimization runs and sufficient for homeostasis. As expected, the strongest effects (bold in Figure 3b) link input chemicals to proliferation regulation, in a mirrored manner across cell types. In both cell types, chemical secreted by the same type inhibits division, while chemical secreted from the other type promotes it. Secretion control in both cell types is governed solely by chemical factor 1, associated with the initially less abundant red cell type. In cell type 1, chemical 1 simply enhances its own production. In cell type 2, secretion regulation is less straightforward: chemical 1 inhibits secretion of chemical 2, and high secretion of chemical 2 contribute to proliferation suppression.
Consider a region where type 2 cells are surrounded by their own type, levels of chemical 1 are low and chemical 2 is secreted at a basal rate. In this case, the lack of chemical 1 removes suppression of chemical 2 secretion, resulting in higher chemical 2 levels. High levels of chemical 2 and strong secretion activate division inhibitory pathways for type 2, halting proliferation. Conversely, when sensed levels of chemical 1 rise due to increased division of cell type 1, the inhibition of chemical 2 secretion is lifted, eliminating in turn the secondary inhibitory feedback on proliferation. Simultaneously, the strong positive regulatory connection between chemical 1 and division promotes type 2 proliferation. Increased type 2 division and secretion of chemical 2 again activates the inhibitory division pathways, and the cycle repeats. Notably, spatial structure plays a key role: cell divisions place daughter cells near parent cells, altering neighborhood cell type composition. This increases the local concentration of the self-secreted chemical and, at the same time, decreases the concentration of the other chemical as cells move away from surrounding sources. Despite model differences, the learned solutions share key features with the macrophage-fibroblast homeostasis circuit reported in [39]. Both systems feature mutual proliferation promotion between the two cell types and negative growth self-regulation. Furthermore, also in this case one of the two factors negatively regulate the other.
Beyond the most common networks, optimization also identifies two alternatives for cell type 1 and one for cell type 2. These variants differ from the most frequent architecture by single additional edges (colored differently in Figure 3b). Network variants are not correlated between cell types and can appear in any of the six possible combinations. The weaker additional links enhance existing dynamical behaviors without fundamentally altering network function. Their presence may reflect optimization convergence to local minima or variations in convergence speed, rather than specific functional advantages.
Remarkably, the learned gene regulations exhibit robustness and generalization beyond the explicit optimization goals. If one fixes the initial imbalance to the one used during training ($n_{r}/N=0.2$), the learned regulation maintains homeostasis even in extended simulations far beyond the training duration of 100 divisions, showing no degradation (Fig. 3d). Furthermore, maintaining 100 cell divisions but varying the initial proportions of cell types, homeostasis is reasonably maintained up to nearly twice the initial imbalance (Fig. 3e). Beyond this point, the regulatory circuit starts to overcompensate and loses its advantage over random cell divisions. This behavior aligns with the regulatory asymmetry observed in the discovered networks, reflecting the fact that all initial simulation states during optimization are imbalanced in the same direction (i.e. with more blue cells than red ones). We suppose that the initial tissue state results from an upstream cell differentiation process which does not produce completely random cell type proportions, so having no initial variability seems like a reasonable approximating assumption.
Figure 4: Mechano-chemical Regulation of Homogeneous Growth. (a) Scheme of the problem setup, inspired by wing disc growth in Drosophila. The extrinsic division scaling mirrors the experimentally observed tendency of cells to divide in the center of the cluster. (b) The learned gene network. Division is controlled both by the external division field (left implicit in the diagram) and by the other cell inputs. Edge boldness is correlated to connection strength. (c) Proliferation regulation learned in a power-law division field yield lower coefficient of variation of division propensities also in systems with different proliferation biases. Left to right: the power-law field used in training, a field with two peaks instead of one, an exponential field, and a field with an off-center peak. Distributions are calculated over 100 simulations. (d) Ablation studies assess the importance of both mechanical and chemical regulation. Separately removing chemical or mechanical feedback leads to more proliferation heterogeneity with respect to the presence of both.(e) Division rate divided by the extrinsic division field is plotted for each kind of simulation in panel (c). This shows that cells regulate division probabilities by adapting to the shape of the extrinsic division field.
@@
Mechanical signaling plays a crucial role in morphogenesis, influencing both individual cells and tissue-level processes. At the cellular level, mechanosensation affects gene expression through various biochemical pathways that transduce forces [42]. On a broader scale, mechanical signals convey local instructions to shape global structures [42, 43, 44, 45]. In tissues, mechanical interactions can operate over larger distances than biochemical signals due to tension propagation through connected cells [45]. Furthermore, the accumulation of local mechanical stress can impact the global growth of developing structures. For instance, mechanical feedback has been proposed as a mechanism for growth arrest in the Drosophila wing imaginal disc [46, 47]. In this study, we explore these concepts by optimizing a model inspired by the Drosophila wing disc development to uncover mechanisms that couple mechanics to growth.
Our model incorporates a constant radial growth factor gradient that modulates division propensity, with cells closer to the cluster center having a higher probability of division (Fig. 4a). This aligns with experimental observations of transiently faster growth rates in central cells during early wing disc development [48]. This extrinsic division scaling, fixed throughout the simulation, follows a power-law decay. The total division propensity of cell $i$ is given by:
$d_{i}=g_{i}(stress,chemicals)*f(r)$
where $g_{i}(stress,chemicals)$ is the division output of the internal gene network and $f(r)\propto 1/r^{2}$ is the radial scaling due to the external factor.
As the cell cluster expands, each cell modulates its division propensity through their internal gene circuit, receiving sensed chemicals and stress as inputs. We aim to identify local growth rules that achieve homogeneous proliferation throughout the cluster despite the nonuniform effect of the external factor gradient. This optimization for uniform growth rates is inspired by experimental evidence, showing similar growth rates across different regions of the wing disc during most of its development, despite small-scale local variations [48].
We optimize couplings between genes in the circuit to minimize the coefficient of variation of division propensities: